Can anyone solve this for me.
If 7x y = 23 and x 7y =47, then find the value of x-y?

Found 2 solutions by Alan3354, chidex:
Answer by Alan3354(69443)

Multiply one of the eqns by 7, then subtract.

Answer by chidex(1)

7x y=23-------(1)
x 7y=47-------(2)
from equation (1)
7x y=23
substract 7x from both side
7x-7x y=23-7x
y=23-7x-------(3)
substitu y=23-7x into equation (2)
x 7(23-7x)=47
x 161-49x=47
x-49x=47-161
48x=114
48x/48=114/48
x=2.38 or x=2.4
substitute x=2.4 in equation (2)
7(2.4) y=23
16.4 y=23
y=23-16.4
y=6.6